OpenStudy (anonymous):

One plane at 520 ft. is ascending at a rate of 40 ft. per minute, while another plane at 3800 ft. is descending at a rate of 120 ft. per minute. How long will it take the 2 planes to be at the same altitude?

OpenStudy (jamesj):

You want to write down an equation for the height of both airplanes at time t. Call the height of the first plane a(t); the second plane b(t) Now you know that a(0) = 520 ft a(1) = 520 - 40 = 480 a(2) = 520 - 80 = 440 Hence what's the equation for a(t) in general?

OpenStudy (anonymous):

a(t) - b(t) = 3800 ?

OpenStudy (anonymous):

you add 40 for every time you subtract. like minus 40, minus 80, and now it's going to be minus 120?

OpenStudy (jamesj):

Forget that for moment. What will want to do is solve the equation a(t) = b(t) because that's when the planes are at the same height. But before we get to that equation, we're going to know what a(t) and b(t) is. So look again at the pattern for a(t) and see if you write down the general form of the equation for a(t): And apologies ... I should have added as you say!! a(0) = 520 ft a(1) = 520 + 40 = 560 a(2) = 520 + 80 = 600 a(3) = 520 + 120 = 640

OpenStudy (jamesj):

Hence a(t) = 520 + 40t makes sense?

OpenStudy (anonymous):

and what do i put for a(t) ? 520?

OpenStudy (jamesj):

In general, at t minutes, the first plane is at an altitude of a(t) = 520 + 40t

OpenStudy (jamesj):

So at t = 0 minutes, the plane is where it began, a(0) = 520 At t = 1 minute, it's 40 feet higher a(1) = 520 + 40 = 560 At t = 2 minutes, it's another 40 feet higher a(2) = 520 + 2(40) = 520 + 80 = 600 etc. Make sense?

OpenStudy (jamesj):

Again, the formula for a(t) is a(t) = 520 + 40t

OpenStudy (anonymous):

oh okay! thank you!! :)

OpenStudy (jamesj):

Now, " another plane at 3800 ft. is descending at a rate of 120 ft. per minute. " So what's the equation for b(t)?

OpenStudy (anonymous):

& i have last word problem, if you have time?

OpenStudy (jamesj):

b(0) = 3800 b(1) = 3800 - 120 b(2) = 3800 - 240 ...

OpenStudy (anonymous):

b(t) = 3800 - 120x

OpenStudy (jamesj):

Right. Now solve a(t) = b(t)

OpenStudy (anonymous):

120t *

OpenStudy (anonymous):

okay, now i can do this part by myself:) thank you!

OpenStudy (jamesj):

ok. I've got to go. Good luck.

OpenStudy (anonymous):

:)